In PRS, segment PT is an altitude and segment PX is a median.
Find RS if RX = x + 7 and SX = 3x – 11
Find RT if RT = x – 6 and PTR = 8x - 6

Answer by ikleyn(52786)

In PRS, segment PT is an altitude and segment PX is a median.
Find RS if RX = x + 7 and SX = 3x – 11
Find RT if RT = x – 6 and PTR = 8x - 6
~~~~~~~~~~~~~~~~~~~

1.  Find  RS  if  RX = x + 7  and  SX = 3x – 11

a)  Make a sketch to follow my arguments.


b)  Since PX is the median, |RT| = |SX|  (the measures of these segments are the same).

    It gives you an equation'

    x + 7 = 3x - 11  ====>  7 + 11 = 3x - x  ====>  2x = 18  ====>  x = 9.


    Then |RS| = x+7 = 9+7 = 16;  |SX| = 3x-11 = 3*9-11 = 16  (it is not amazing: it should be so !)

         and  |RS| = 16 + 16 = 32.


2.  Find  RT  if RT = x – 6  and  PTR = 8x - 6.

    Honestly, I can not read this:  PTR = 8x - 6,

    and do not understand what does it mean.


In future posts, please do not use this symbol "  ", since it is non-standard and its meaning is dark.